OpenStudy (anonymous):

Simplify: t^2-25/t^2+8t+15/t-5/t+9

OpenStudy (anonymous):

(t^2-25)(t+9) / (t^2+8t+15)(t-5)

OpenStudy (anonymous):

how do you input fractions?

OpenStudy (anonymous):

use \frac{}{} (What I put is same as what u put)

OpenStudy (anonymous):

It's a fraction divided by a fraction, right?

OpenStudy (anonymous):

Yes, so flip the bottom fraction and multiply instead... (t^2-25)(t+9) / (t^2+8t+15)(t-5)

OpenStudy (anonymous):

\[\frac{t ^{2}}{t ^{2}+8t+15}/\frac{t-5}{t+9} \] turns in to \[\frac{(t ^{2}-25)(t+9)}{(t ^{2}+8t+15)(t-5)}\]

OpenStudy (aroub):

(t-5)(t+5)(t+9)/ (t+5)(t+3) (t-5) (t+9)/(t+3)

OpenStudy (aroub):

hey ajmal!! :D

OpenStudy (anonymous):

hey aroub!!!!!:)

OpenStudy (aroub):

(t+9)/(t+3) is the answer..

OpenStudy (aroub):

ok after factoring.. u simplify the same numbers in the numerator and the denominator meaning for example in here u have (t+5)(t-5) in the numerator and the denominator so u cancel them out..

OpenStudy (anonymous):

You all are geniuses... Wish I had your brains for math.